### v1

@@ 8 Facts and Language Generation
In such paradigms, one natural place where quantum computing has potential to enhance performance is in accessing and designing the knowledge base. Grover’s algorithm (Grover, 1996) yields a quadratic speedup over classical methods for searching a database, and hence forms the backbone of quantum information retrieval systems (Giri and Korepin, 2017). Going one step further, Pronin and Ostroukh (2023) propose using Grover’s search to inform the design of the database itself, and specifically target vector databases as the intended application. However, they do not directly compare the quadratic speedup compares with the complexities achievable by using efficient classical data structures, and a potential interesting direction in this area is the question of how (if possible) to design a practical hybrid knowledge base that combines the best of both classical and quantum processing.
Although techniques such as retrieval-based and knowledge-based generation are a new area in the present-day context of ‘fixing LLMs’, methods such as that of Borgeaud et al. (2022) hearken back to an older class of designs where the ‘facts’ are stored in a knowledge base, and the language model is effectively a source of templates, not of facts. For example, in Figure 19, a knowledge base is used to find the variable that satisfies the question “When was J.S. Bach born?” (the answer being ‘1685’), and then a language model is used to express this as the sentence “J.S. Bach was born in the year 1685.” (A standard early use of such designs was in ‘mail merging’, where a template for a message is combined with a list of different names to generate personalized messages.) Such a language model can just as easily generate the sentence “J.S. Bach was born in 1985”, not because it’s ‘hallucinating’, but because it’s working!
More generally, probabilistic language models are designed to note that Wednesday and Thursday are similar, and so having seen the phrase “Let’s meet on Wednesday”, the model should judge the phrase “Let’s meet on Thursday” to be similarly plausible. Saying that such a probabilistic model ‘hallucinates’ when it generates an untrue sentence reflects a fundamental misunderstanding of probabilistic models. Sampling from a probabilistic model is like rolling dice: if we previously observe a 3 and the dice-roll gives a 4, the dice aren’t ‘hallucinating’ a 4 instead of the ‘true’ value of 3. The problem lies in the assumption that plausible probabilistic samples of language should correspond to facts at all.
@@
The practice of generating text from just a language model was popularized by successful machine translation systems (Bahdanau et al., 2014). With machine translation, it makes sense that the system is not responsible for factual accuracy, because this is the user’s responsibility. In concrete terms, a correct translation of “J.S. Bach was born in 1985” from English to German might be “J.S. Bach wurde 1985 geboren”, not “Input error: prompt contains factual inaccuracy.” Gradually models such as GPT (Floridi and Chiriatti, 2020) demonstrated that a whole range of prompts, not just translation targets, could elicit plausible and fluent outputs. The Chomskian program claimed grammatical fluency as the heart of language decades ago: today, we are seeing that this fluency is another aspect of human behavior that computers can mimic effectively, and that the ability to gabble fluently is only part of the intelligence puzzle.
@@
Quantum theory intersects with these topics even more fundamentally, by explicitly distinguishing the possible from the actual. A quantum circuit has many possible outcomes that could be observed, but only one outcome will be observed when measured: and this fixes the hypothetical situation so that the same outcome will be observed next time. A multiplicity of possibilities can become a single fixed event (Heisenberg, 1958). Formal similarities between this process and language ambiguity were noted by Widdows (2003), and the quantum economic theory of Orrell (2020) is based on the use of quantum information to model beliefs about values, and classical information to model amounts of money agreed in fixed transactions.
@@
The problem of distinguishing things that might happen from things that do happen was behind some of the controversies of early classical mechanics. Leibniz discussed the notion of possible worlds, and maintained that there must be a rational necessity behind (God’s) choosing this world (Rescher, 1996). Newton’s belief in ‘absolute space’ implied a fixed zero-point or ‘origin’, and Leibniz argued that this implied that God must have made an arbitrary choice without a necessary reason, which was unacceptable (Bouquiaux, 2008). Such considerations of necessity vs. contingency and their relationship to past, present, and future in time, go back at least to the famous ‘sea-fight’ discussion in Aristotle’s De Interpretatione (McKeon, 1941).
@@
The notion that there are different ‘possible worlds’ where a macroscopic event did or did not happen, that one of those worlds is ‘chosen’ based on a small local decision, and this possible world thus becomes the actual world, was thrust into the limelight by quantum mechanics. The implication of superposition and large-scale randomness was troubling to Einstein (“God does not play dice!”) and Schrödinger, whose famous paradoxical cat was designed to illustrate the absurdity of quantum mechanics in large-scale reality, where “the working of an organism requires exact physical laws” Schrödinger (1944). By contrast, Bohr and Heisenberg supported the ‘Copenhagen Interpretation’, where the wave-function represents real possibilities, and “the transition from the “possible” to the “actual” takes place during the act of observation” (Heisenberg, 1958, Ch 3).
@@
Some of the challenges inherent in large stochastic problems, like weather forecasting, are thus philosophically related to key questions of how one possible future is selected and becomes the past. Quantum mechanics does not completely answer this question, but it does better than classical mechanics, where the assumption of a deterministic universe avoids the problem. Heisenberg’s analysis of where different uncertainties come from, and how we should think about them, has useful insights including “This probability function represents a mixture of two things, partly a fact and partly our knowledge of a fact” (Heisenberg, 1958, Ch 3). This does not tell us how to fix language models, but it is a good reminder that our ways of stating and communicating ‘facts’ are entirely human. Practically, it helps to understand probabilistic language models as generators of hypothetical utterances, rather than factual statements, and the generative nature of language models is precisely what enables them to go smoothly from data they encountered to data they might just as well have encountered. In a sense, a large language generator is a kind of hypothesis-generator with the gift of the gab. Language models do this task very well, but this should never have convinced us that a model will generate truthful language without an independent source of knowledge. With these considerations, quantum theory has some insight on potential solutions to improve language modeling systems, and at least guards against mistakes that arise from over-deploying hypothesis-generation systems without suitable observation processes.

### v2

@@ 8 Facts and Language Generation
In such paradigms, one natural place where quantum computing has potential to enhance performance is in accessing and designing the knowledge base. Grover’s algorithm (Grover, 1996) yields a quadratic speedup over classical methods for searching a database, and hence forms the backbone of quantum information retrieval systems (Giri and Korepin, 2017). Going one step further, Pronin and Ostroukh (2023) propose using Grover’s search to inform the design of the database itself, and specifically target vector databases as the intended application. However, they do not directly compare the quadratic speedup with the complexities achievable by using efficient classical data structures, and a potential interesting direction in this area is the question of how (if possible) to design a practical hybrid knowledge base that combines the best of both classical and quantum processing.
Although techniques such as retrieval-based and knowledge-based generation are a new area in the present-day context of fixing LLMs, methods such as that of Borgeaud et al. (2022) hearken back to an older class of designs where the facts are stored in a knowledge base, and the language model is effectively a source of templates, not of facts. The spotlight on language generation in the past few years has refocused work on such methods, and how best to combine them with LLMs Zhang et al. (2023). For example, in Figure 21, a knowledge base is used to find the variable that satisfies the question “When was J.S. Bach born?” (the answer being “1685”), and then a language model is used to express this as the sentence “J.S. Bach was born in the year 1685.” (A standard early use of such designs was in mail merging, where a template for a message is combined with a list of different names to generate personalized messages.) Such a language model can just as easily generate the sentence “J.S. Bach was born in 1985”, not because it’s hallucinating, but because it’s working correctly with a different knowledge base.
More generally, probabilistic language models are designed to note that Wednesday and Thursday are similar, and so having seen the phrase “Let’s meet on Wednesday”, the model should judge the phrase “Let’s meet on Thursday” to be similarly plausible. Saying that such a probabilistic model “hallucinates” when it generates an untrue sentence reflects a fundamental misunderstanding of probabilistic models. Sampling from a probabilistic model is like rolling dice: if we previously observe a 3 and the dice-roll gives a 4, the dice aren’t hallucinating a 4 instead of the “true” value of 3. The problem lies in the assumption that plausible probabilistic samples of language should correspond to facts at all.
@@
The practice of generating text from just a language model was popularized by successful machine translation systems (Bahdanau et al., 2014). With machine translation, it makes sense that the system is not responsible for factual accuracy, because this is the user’s responsibility. In concrete terms, a correct translation of “J.S. Bach was born in 1985” from English to German might be “J.S. Bach wurde 1985 geboren”, not “Input error: prompt contains factual inaccuracy.” Gradually models such as GPT demonstrated that a whole range of prompts, not just translation targets, could elicit plausible and fluent responses (Brown et al., 2020). The Chomskian program claimed grammatical fluency as the heart of language decades ago: today, we are seeing that this fluency is another aspect of human behavior that computers can mimic effectively; and the ability to assemble erudite text has become one of the most impressively-solved parts of AI, sometimes leading to problems elsewhere.
@@
Quantum theory intersects with these topics even more fundamentally, by explicitly distinguishing the possible from the actual. A quantum circuit has many possible outcomes that could be observed, but only one outcome is observed when measured: and this fixes the hypothetical situation so that the same outcome is observed next time. A multiplicity of possibilities can become a single fixed event (Heisenberg, 1958). Formal similarities between this process and language ambiguity were noted by Widdows (2003), and the quantum economic theory of Orrell (2020) is based on the use of quantum information to model beliefs about values, and classical information to model amounts of money agreed in fixed transactions.
@@
The problem of distinguishing things that might happen from things that do happen was behind some of the controversies of early classical mechanics. Leibniz discussed the notion of possible worlds, and maintained that there must be a rational necessity behind (God’s) choosing this world (Rescher, 1996). Newton’s belief in absolute space implied a fixed zero-point or origin, and Leibniz argued that this implied that God must have made an arbitrary choice without a necessary reason, which was unacceptable (Bouquiaux, 2008). Such considerations of necessity vs. contingency and their relationship to past, present, and future in time, go back at least to the famous sea-fight discussion in Aristotle’s De Interpretatione (McKeon, 1941).
@@
The notion that there are different possible worlds where a macroscopic event did or did not happen, that one of those worlds is chosen based on a small local decision, and this possible world thus becomes the actual world, was thrust into the limelight by quantum mechanics. The implication of superposition and large-scale randomness was troubling to Einstein (“God does not play dice!”) and Schrödinger, whose famous paradoxical cat was designed to illustrate the absurdity of quantum mechanics in large-scale reality, where “the working of an organism requires exact physical laws” Schrödinger (1944). By contrast, Bohr and Heisenberg supported the Copenhagen Interpretation, where the wave-function represents real possibilities, and “the transition from the “possible” to the “actual” takes place during the act of observation” (Heisenberg, 1958, Ch 3).
@@
Some of the challenges inherent in large stochastic problems, like weather forecasting, are thus philosophically related to key questions of how one possible future is selected and becomes the past. Quantum mechanics does not completely answer this question, but it does better than classical mechanics, where the assumption of a deterministic universe avoids the problem. Heisenberg’s analysis of where different uncertainties come from, and how we should think about them, has useful insights including “This probability function represents a mixture of two things, partly a fact and partly our knowledge of a fact” (Heisenberg, 1958, Ch 3). This does not tell us how to fix language models, but it is a good reminder that our ways of stating and communicating facts are entirely human. Practically, it helps to understand probabilistic language models as generators of hypothetical utterances, rather than factual statements, and the generative nature of language models is precisely what enables them to go smoothly from data they encountered to data they might just as well have encountered. In a sense, a large language generator is a kind of hypothesis-generator with the gift of the gab. Language models do this task very well, but this should never have convinced us that a model will generate truthful language without an independent source of knowledge. With these considerations, quantum theory has some insight on potential solutions to improve language modeling systems, and at least guards against mistakes that arise from over-deploying hypothesis-generation systems without suitable observation processes.